How to go back in a table so the common difference is adding each time by 5,7,9 etc
1
Expert's answer
2019-08-29T08:27:24-0400

Let us start a sequence from "f(0)=a". We have next table:



"\begin{matrix}
x & 0 & 1 & 2 & \dots & n \\
f(x) & a & a+5 & a+5+7 & \dots & a+\sum\limits_{k=1}^n (2k+3)
\end{matrix}"


So, "f(n)-f(n-1)=a+\sum\limits_{k=1}^{n} (2k+3)-a-\sum\limits_{k=1}^{n-1} (2k+3)=2n+3" and "2n+3 \in \{5,7,9,\dots\}"

Moreover, "f(n)=a+\sum\limits_{k=1}^n (2k+3)=a+2\frac{n(n+1)}{2}+3n=a+n^2+4n"